OpenStudy (camerondoherty):

A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

Okay so in order to do this you want to multiply the second line by -4

OpenStudy (camerondoherty):

ok so how?

OpenStudy (anonymous):

Multiply every term in the second like by (-4)

OpenStudy (camerondoherty):

oh kk

OpenStudy (anonymous):

(-4)(2x+2y)= 3(-4)

OpenStudy (camerondoherty):

-8x + -8y = -12

OpenStudy (anonymous):

So now you can just cancel out the y if you subtract the equations. 3x + 8y = 12 -8x -8y = -12

OpenStudy (anonymous):

So you would get -5x=0 then x=0, and to solve for y plug in x into any of the two.

OpenStudy (anonymous):

Y=2/3

OpenStudy (camerondoherty):

wait so y=0?

OpenStudy (anonymous):

Wait no y=3/2

OpenStudy (camerondoherty):

so x=0 then?

OpenStudy (anonymous):

You just plug in x into any of the two equations

OpenStudy (anonymous):

Yes that's what we solved first

OpenStudy (camerondoherty):

ok so x=0 and y=3/2 and you substitute the values and solve

OpenStudy (anonymous):

3(0) + 8y=12 then 8y=12 y=12/8 or 3/2

OpenStudy (camerondoherty):

to check

OpenStudy (anonymous):

Yeah

OpenStudy (camerondoherty):

oh ok than i have to solve the original right?

OpenStudy (anonymous):

We did solve the original. But umm I don't really get what the directions is asking for.

OpenStudy (camerondoherty):

lol ok

OpenStudy (anonymous):

3x+8y=12 -8x-8y=-12 would be an equivalent system and they would give you the same answer

OpenStudy (camerondoherty):

Thanks a bunch!

OpenStudy (anonymous):

Yeah, no problem babe.

OpenStudy (anonymous):

3x + 8y = 12 2x + 2y = 3 -6x = 16y = -24 6x + 6y = 9 -10y = -15 /-10 /-10 y = 1.5 sub in y into the equation and solve for x :) @camerondoherty

OpenStudy (camerondoherty):

You copied pasted that -_- http://prntscr.com/41o8av

OpenStudy (anonymous):

yeah... Atleast I helped you

OpenStudy (anonymous):

Anyways yeah you get the same answer y=3/1 or 1.5

OpenStudy (anonymous):

I don't think shes online

OpenStudy (anonymous):

Monster1234 take your plagiarism and leave this site. we don't want help from cheaters.

OpenStudy (anonymous):

It's Russian

OpenStudy (anonymous):

You could have linked that question instead of submitting it as your own

OpenStudy (anonymous):

atleast you people actually hep others right? so why be so ungrateful